factor the expression completely
x^2 + 10x + 25
I need a step by step process to follow for quadratic expression factoring. I cant get the answer correct.I have tried over and over.
I have tried (1x+5)(1x+5) and (x+5)(x+5)???

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

factor the expression completely
x^2 + 10x + 25
I need a step by step process to follow for quadratic expression factoring. I cant get the answer correct.I have tried over and over.
I have tried (1x+5)(1x+5) and (x+5)(x+5)???
-----------------
I have tried (1x+5)(1x+5) and (x+5)(x+5)???
Those are the same, and that's the answer.

Answer by ewatrrr(24785)

 
Hi,
(x+5)(x+5) = x^2 + 10x + 25 = (x+5)(x+5)
F First terms x^2
O Outside terms 5x
I Inside terms 5x
L Last terms 25
Outside terms + Inside terms = 5x + 5x = 10x